Three uncharged capacitors with equal capacitances are combined in series. The combination is connected to a 7.35-V battery, which charges the capacitors. The charging process involves 2.63 × 10-4 C of charge moving through the battery. Find the capacitance of each capacitor.
Answer #1

We want to solve for C, the capacitance of one of the capacitors.
The equivalent capacitance of the whole system is 1/(1/C + 1/C + 1/C) = 1/(3/C) = C/3
Capacitance has units of Farads or Coulombs/Volt so all we need to do is divide the amount of charge that moves through the battery during the charging process by the voltage across the battery, and that will give the total capacitance of the system:
(2.63 x 10^-4 C) / (7.35 V) = 3.58 x 10^-5 C
Now we have the total capacitance of the system, which is equal to C/3, so to solve for C:
(3.58 x 10^-5 C) = C/3
C = 1.074 x 10^-4 F